# Advancements in Constitutive Model Calibration: Leveraging the Power of Full-Field DIC Measurements and In-Situ Load Path Selection for Reliable Parameter Inference

### v1

@@ 4.2 Results
Model calibration was first performed with all possible load paths with an initial load step along Axis A for a total of 16 calibrations with pre-determined load paths. For ease of reading, A is used to denote a load step along Axis A (horizontal direction) and B for Axis B (vertical direction). Table 4 reports the expected values, $\mathbb{E}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}$, marginal variances, $\mathbb{V}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}$ and 95% credible intervals (CI) for parameter vector $\boldsymbol{\theta}=[\sigma_{y},A,n,a]$ for all 16 calibrations, with $\sigma_{y}$ and $A$ being reported in units of ksi. The marginal variance is the variance (or uncertainty) of one variable when looked at in isolation, without consideration of the others, and the 95% CI is the interval in which the parameter has a 95% probability to fall within. Since a Laplace approximation to the posterior distribution was used, the marginal posterior distributions are symmetric, and the 95% CIs are reported with a single value, $\zeta$, where the expected value plus and minus $\zeta$ defines the interval: $95\%\,\text{CI}=\mathbb{E}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}\,\pm\zeta$. Additionally, the collective parameter uncertainty, reported as the determinant of the posterior covariance $\det(\Sigma)$ (a.k.a. the generalized variance), is included.
The resultant parameter expected values were very similar for each load path. Specifically, $\mathbb{E}_{\sigma_{y}\>|\>\boldsymbol{y},\boldsymbol{\xi}}$ for the initial yield stress was in agreement to the tenth for all cases, and the expected value of all other parameters were in close agreement to the tenth. The marginal variances for $\sigma_{y}$ and $a$ were similar for all cases; however, there were orders of magnitude difference in the marginal variances reported for $A$ and $n$ among the different cases (although still on a relatively small scale). There was likewise orders of magnitude difference in the reported generalized variance among the different load paths. A similar trend was seen in the 95% CI; the interval was similar for parameters $\sigma_{y}$ and $a$ for all cases, but showed more deviation case to case for parameters $A$ and $n$. Thus parameters $A$ and $n$ were the most sensitive to the applied load path and resulting calibration data. In summary, the pre-determined AAAAA load path exhibited the least amount of parameter uncertainty, expressed through the various reported uncertainty metrics, thus demonstrating the highest confidence in the parameter values; likewise, the ABABA load path had the greatest parameter uncertainty.
@@
AAAAB [42.5, 13.6, 14.4, 11.1] [ 2.0, 17.9, 84.7, 4.9 ] [0.03, 0.08, 0.18, 0.04] 1.8
AAABA [42.5, 13.6, 14.3, 11.2] [ 1.9, 24.5, 112.5, 5.2 ] [0.03, 0.10, 0.21, 0.04] 3.7
AAABB [42.5, 13.7, 14.2, 11.1] [ 1.7, 50.2, 143.2, 6.2 ] [0.03, 0.14, 0.23, 0.05] 10.0
AABAA [42.5, 13.5, 14.6, 11.2] [ 1.8, 15.1, 87.3, 6.3 ] [0.03, 0.08, 0.18, 0.05] 4.5
AABAB [42.5, 13.5, 14.5, 11.2] [ 1.8, 94.8, 299.6, 6.0 ] [0.03, 0.19, 0.34, 0.05] 35.1
AABBA [42.5, 13.6, 14.4, 11.1] [ 3.0, 63.8, 238.6, 7.9 ] [0.03, 0.16, 0.30, 0.06] 51.9
AABBB [42.5, 13.7, 14.3, 11.2] [ 2.5, 14.4, 96.4, 7.3 ] [0.03, 0.07, 0.19, 0.05] 6.8
ABAAA [42.5, 13.6, 14.4, 11.3] [ 1.4, 11.7, 69.3, 11.2 ] [0.02, 0.07, 0.16, 0.07] 5.4
ABAAB [42.5, 13.3, 14.9, 11.2] [ 2.3, 114.4, 484.6, 8.4 ] [0.03, 0.21, 0.43, 0.06] 55.5
@@
ABABB [42.5, 13.5, 14.7, 11.2] [ 1.9, 26.6, 131.0, 13.0 ] [0.03, 0.10, 0.22, 0.07] 48.3
ABBAA [42.5, 13.6, 14.5, 11.2] [ 1.4, 31.7, 96.9, 10.8 ] [0.02, 0.11, 0.19, 0.06] 29.0
ABBAB [42.5, 13.8, 14.2, 11.2] [ 1.7, 168.6, 442.2, 7.1 ] [0.03, 0.25, 0.41, 0.05] 92.9
ABBBA [42.5, 13.6, 14.4, 11.2] [ 1.9, 108.6, 379.1, 7.5 ] [0.03, 0.20, 0.38, 0.05] 29.0
ABBBB [42.5, 13.6, 14.3, 11.2] [ 1.3, 8.1, 47.6, 9.9 ] [0.02, 0.06, 0.14, 0.06] 2.49
@@
Table 4: Posterior summaries for 16 pre-determined load paths, each starting within an initial A load step. The parameter expected values, $\mathbb{E}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}$, marginal variances, $\mathbb{V}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}$, and 95% CI are reported for parameter vector $\boldsymbol{\theta}=[\sigma_{y},A,n,a]$ ($\sigma_{y}$ and $A$ reported in units of ksi) along with the generalized variance, which is the determinant of the posterior covariance. The 95% CI is represented as $\mathbb{E}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}\,\pm\zeta$, where $\zeta$ is the reported value. For visual simplicity, horizontal load steps are reported with an A and vertical load steps with a B. The material model parameters used for data generation were $\boldsymbol{\theta}^{true}=[42.51,13.63,14.35,11.19]$.
@@
In reality, all 16 possible load paths would not be evaluated experimentally, since the goal of this work is to determine the one optimal path. Therefore, the load path was then chosen adaptively, as would be done in reality, with the ICC frame work by calculating the EIG of the candidate load steps in $\Xi$ at each decision point as the cruciform was actively guided through the load path tree. The EIG of each candidate load step is shown in Fig. 9 for each decision point. The EIG steadily decreases as the load steps progress, implying that additional data—for either load step option—has a waning influence on the calibration. An A load step was chosen each time for a total load path of AAAAA. As was confirmed with the pre-determined load paths in Table 4, this is in fact the load path with the least collective parameter uncertainty given the data, so while it was not guaranteed, in this case the myopic decision-making did result in the globally optimal load path (given the preset tree structure).
Figure 9: The EIG for each candidate load step in the load path.
Fig. 10 shows the posterior contours for the AAAAA load path (a), as well as a side-by-side comparison with contours for an ABABA load path (b) so that a visual comparison can be made between the cases with least and greatest parameter uncertainty. Contours for the ABABA load path are shown separately in Fig. 3 in G. The diagonal plots in (a) and (b) show the marginal posterior distribution for each parameter along with the 95% credible interval (CI) (dashed black lines) as well as $\boldsymbol{\theta}^{true}$ (pink dashed line). The off-diagonal plots show the joint marginal posterior contours for each parameter pair along with $\boldsymbol{\theta}^{true}$ indicated by a pink star. The parameter uncertainty is described by the posterior distribution, and the MAP parameter estimate corresponds with the value which maximizes the distribution.
The two load paths resulted in posterior distributions that exhibited different marginal and joint parameter uncertainty, as well as different parameter correlations (reported in E). This visual comparison makes it clear that even though both load paths resulted in expected values that were very similar (Table 4), the results of the inference are not the same. Importantly, load path AAAAA demonstrated lower parameter uncertainty compared to load path ABABA. The selection of an AAAAA load path over ABABA, or any other alternative, is not necessarily intuitive a priori. This result shows that the BOED algorithm was successfully able to identify the load path that produced more informative data than an alternative load path which may be selected a priori by subject-matter expertise.
[…]
[figure]style=plain,subcapbesideposition=top,font=small
@@
Figure 10: The posterior contour for load path (a) AAAAA and (b) both AAAAA and ABABA. Marginal posterior densities are plotted on the diagonals and joint-marginal densities on the off-diagonals. $\boldsymbol{\theta}^{true}$ is indicated with a dashed pink line and star and the 95% CI with dashed black lines.
The posterior probability is visualised over the surrogate model output (load and PCA singular values) by drawing 200 parameter samples from the posterior distribution $\boldsymbol{\theta}^{draw}\sim\pi(\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi})$ and evaluating the surrogate model at each sample $\tilde{g}(\boldsymbol{\theta}^{draw})$. The corresponding posterior probability of the load values at load step 5 is shown in Fig. 11 and for the PCA singular values in Fig. 3 in F for load path AAAAA. The model output at the MAP parameter estimate is shown with a dotted black line, and the 95% CI of the distribution over the load is shown with dashed black lines, indicating the interval in which the load values have a 95% probability of residing within.
The singular values obtained from $\tilde{g}(\boldsymbol{\theta}^{draw})$ were then used along with the 5 retained principal components to reconstruct the displacement field, $\text{disp}_{i}^{\text{recon}}=\text{disp}^{\text{PCA}}_{i}V^{*T}_{i}+\text{disp}^{\text{mean}}_{i},\,i\in[X,Y]$, so that there were 200 reconstructed instances of the displacement field. Fig. 12 shows the observed displacement field generated with $\boldsymbol{\theta}^{true}$ at the final load step for the AAAAA load path along with two line scans, across which the 95% CI was computed from the 200 reconstructed fields and plotted in Fig. 13. Given the relative scale of the displacements and the low uncertainty, the 95% CI interval appears as a single line. Overall, the high confidence in the parameter values translates to a high confidence in the resulting model output.
Figure 11: Posterior probability over the load for the AAAAA load path for both the $X$ (left column) and $Y$ (right column) directional components. The load at the MAP parameter estimate is shown with a dotted black line, and the 95% CI of the distribution over the load is shown with dashed black lines, indicating the interval in which the load values have a 95% probability of residing within.
[…]
Figure 12: The observed displacement field generated with $\boldsymbol{\theta}^{true}$ and added noise for the (a) X direction and (b) Y direction of the AAAAA load path at the final load step reported in microns. Lines (1) and (2) indicate the space over which the posterior probability of the displacement field is plotted in Fig. 13.
@@
Figure 13: Posterior probability over the nodal displacements for the AAAAA load path across the two lines indicated in Fig. 12. The $X$ (left column) and $Y$ (right column) directional components are shown along with Line 1 (top row) and Line 2 (bottom row). The 95% CI of the distribution over the displacement is shown with dashed black lines. Given the relative scale of the displacements and the low uncertainty, the 95% CI interval appears as a single line. The x-axis is the normalized location across each line scan.
Results for the ABABA load path calibration, which yielded the highest parameter uncertainty of the pre-determined load paths—Table 4—can be found in G. The observed data (Figs.1 & 2), along with the posterior uncertainty over the load (Fig. 4), PCA singular values (Fig. 5), and displacement field line scans (Figs. 6 & 7) are included. Similar results to the AAAAA case were observed: The posterior probability over the parameters translated to high confidence in the model output and recovered displacement field.
@@ 4.3 Model Validation
Figure 14 shows the posterior predictive distribution over the load for an ABBBA load path for the AAAAA-calibrated model. The distribution over the load is shown for each load step and for each directional component, $X$ and $Y$. Also shown is the observed data (solid red line) generated with $\boldsymbol{\theta}^{true}$ and added noise for the new load path, the predicted load at the MAP parameter estimate (dotted black line), and the 95% CI (dashed black lines). For each load step and for both directional components, the observed data is well within the 95% CI and has high probability under the assumed model.
@@
Figure 14: The posterior predictive distribution for an ABBBA load path is shown over the load for the $X$ (left column) and $Y$ (right column) directional components and at each load step. The observed data is shown with a solid red line, the load at the MAP parameter estimate is shown with a dotted black line, and the 95% CI is shown with dashed black lines.
@@
Visualizing the posterior predictive distribution over the displacement field is more challenging; two complementary visualizations are utilized here. Although it is challenging to visualize the uncertainty over the entire field, 3 of the 200 samples from the posterior predictive distribution were used to reconstruct corresponding displacement fields and are visualized in Fig. 15 for the X directional component. The observed data obtained at the end of each of the 5 load steps of the ABBBA loadpath is shown along with 3 samples from the predictive distribution. This allows for a comparison to be made between the displacement field that is observed and that which is expected to be observed under the posterior distribution that was obtained. The multiple predictive samples serve to establish repeatability, or that the patterns seen in the displacement field did not merely happen by chance. There is no visible systematic discrepancy between the observed and predicted data, and it is reasonable to suggest that the observed data came from the same distribution as the predictions. Similar results were seen for the Y directional component—Fig. 4 in F.
@@
Figure 15: Samples from the posterior predictive distribution of an AAAAA load path is shown for the $X$ directional component of the displacement field for a new ABBBA load path. The left column shows the observed data for each of the 5 load steps (LS) and the three right columns show three different predictive samples.
@@
The 95% CI of the distribution was calculated over the field point-wise from the 200 samples. Fig. 16 shows a full-field contour plot from which the 95% CI over Line 1 was extracted and plotted in Fig. 17 for all 5 load steps (LS) to show the predictive uncertainty. Plots in the left column show the 95% CI (dashed lines) along with the observed data (red line). Since distinguishing between the lines is difficult in some plots, the plots in the right column show the difference between the 95% CI and the observed data. The solid line represents the observations (centered at 0 difference), and the difference between the observations and the CIs are shown with dashed lines. Everywhere the observations fall outside the 95% CI coincides with locations where the upper bound difference is negative or the lower bound difference is positive; these regions are colored orange. Very few points along Line 1 have observations that fall outside the 95% CI. Similar observations were made for the predictive distribution over Line 2 as is shown in Fig. 5 in F.
@@
Figure 16: The observed displacement field for the (a) X and (b) Y directional components of the ABBBA load path at the final load step reported in microns. Lines 1 and 2 mark the space over which the posterior predictive distribution is plotted in Figs. 17 and 5.
Figure 17: The posterior predictive distribution (from an AAAAA load path) over Line 1 displacements (Fig. 16) is shown for new load path ABBBA. The two left columns show the 95% CI of the distribution as well as the observations for the $X$ and $Y$ directional components for all 5 load steps (LS). The two right columns show the difference between the 95% CI and the observations. Regions where the observations fall outside the 95% CI are colored orange, and 0 is marked with a red line. the x-axis is the normalized location along Line 1.
Similar plots were made for the posterior predictive distribution from the ABABA calibration for the new ABBBA load path in G.3 to serve as a visual comparison of the two cases. Additionally, a quantitative comparison is made between the predictive accuracy of the AAAAA vs ABABA calibrations in F.4 through the log predictive density (or log likelihood). This predictive score slightly favored the AAAAA load path overall.
In summary, the model validation showed that the model obtained through the calibration fit the data well. The posterior predictive distribution was visualized over the load at each load step as well as over the displacement field. There were no systematic discrepancies seen between the observed data and the predicted data. Additionally, the observed data looked plausible under the posterior predictive distribution; generally speaking, the observed data had high probability under the assumed model and fell within the 95% CI. The predictive accuracy of the AAAAA- and ABABA-calibrated models was compared visually through plots of the predictive density over the load and displacement field as well as quantitatively through their best-fit log predictive density of 16 new data sets not seen in calibration. The results showed that the AAAAA load path generally yielded greater predictive probability than the ABABA load path. Thus, the well-calibrated and validated model with uncertainty metrics can be used for decision-making with a quantified level of confidence. Importantly, the ICC framework achieved this efficiently and with greater credibility than would have been attained through alternative load paths.

### v3

@@ 4.2 Results
Model calibration was first performed with all possible load paths with an initial load step along the X axis for a total of 16 calibrations with pre-determined load paths. For ease of reading, A is used to denote a load step along the X axis (horizontal direction) and B for a load step along the Y axis (vertical direction). Table 4 reports the expected values, $\mathbb{E}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}$, marginal variances, $\mathbb{V}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}$ and 95% credible intervals (CI) for parameter vector $\boldsymbol{\theta}=[\sigma_{y},A,n,a]$ for all 16 calibrations, with $\sigma_{y}$ and $A$ being reported in units of ksi. The marginal variance is the variance (or uncertainty) of one variable when looked at in isolation, without consideration of the others, and the 95% CI is the interval in which the parameter has a 95% probability to fall within. Since a Laplace approximation to the posterior distribution was used, the marginal posterior distributions are symmetric, and the 95% CIs are reported with a single value, $\zeta$, where the expected value plus and minus $\zeta$ defines the interval: $95\%\,\text{CI}=\mathbb{E}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}\,\pm\zeta$. Additionally, the generalized variance, which is a valuable measure of overall variance in a multivariate distribution, defined as the determinant of the posterior covariance $\det(\Sigma)$, is included. It captures the joint variability of all variables, providing a single scalar value that reflects the spread and orientation of the data. A larger generalized variance indicates greater dispersion among the variables, while a smaller value suggests closer clustering around the mean.
The resultant expected values exhibited less sensitivity to the load path than the uncertainty about their values. Parameter expected values were very similar for each load path. Specifically, $\mathbb{E}_{\sigma_{y}\>|\>\boldsymbol{y},\boldsymbol{\xi}}$ for the initial yield stress was in agreement to the tenth for all cases, and the expected value of all other parameters were in close agreement to the tenth. The marginal variances for $\sigma_{y}$ and $a$ were similar for all cases; however, there were orders of magnitude difference in the marginal variances reported for $A$ and $n$ among the different load paths. There was likewise orders of magnitude difference in the reported generalized variance among the different load paths. A similar trend was seen in the 95% CI; the interval was similar for parameters $\sigma_{y}$ and $a$ for all cases, but showed more deviation case to case for parameters $A$ and $n$. Thus parameters $A$ and $n$ were the most sensitive to the applied load path and resulting calibration data. In summary, the pre-determined AAAAA load path resulted in the least amount of parameter uncertainty (highlighted in green), expressed through the various reported uncertainty metrics, thus demonstrating the highest confidence in the parameter values; conversely, the ABABA resulted in the greatest parameter uncertainty (highlighted in red).
@@
AAAAB [42.51, 13.56, 14.45, 11.17] [ 1.9, 17.4, 84.0, 5.0 ] [0.03, 0.08, 0.18, 0.04] 1.8
AAABA [42.50, 13.46, 14.66, 11.17] [ 1.9, 21.8, 113.7, 5.2 ] [0.03, 0.09, 0.21, 0.04] 3.6
AAABB [42.50, 13.63, 14.32, 11.16] [ 1.8, 49.4, 145.3, 6.3 ] [0.03, 0.14, 0.24, 0.05] 10.1
AABAA [42.51, 13.54, 14.52, 11.19] [ 1.8, 15.8, 89.6, 6.4 ] [0.03, 0.08, 0.19, 0.05] 4.6
AABAB [42.50, 13.53, 14.50, 11.19] [ 1.8, 97.5, 305.9, 6.1 ] [0.03, 0.19, 0.34, 0.05] 35.9
AABBA [42.52, 13.59, 14.46, 11.23] [ 2.9, 61.5, 230.8, 8.1 ] [0.03, 0.15, 0.30, 0.06] 51.4
AABBB [42.50, 13.61, 14.56, 11.23] [ 2.5, 13.7, 96.9, 7.5 ] [0.03, 0.07, 0.19, 0.05] 6.9
ABAAA [42.50, 13.59, 14.39, 11.17] [ 1.4, 11.7, 68.7, 10.8 ] [0.02, 0.07, 0.16, 0.06] 5.2
ABAAB [42.50, 13.54, 14.50, 11.19] [ 2.3, 145.2, 524.9, 8.2 ] [0.03, 0.24, 0.45, 0.06] 61.5
@@
ABABB [42.51, 13.60, 14.44, 11.24] [ 1.9, 28.8, 130.7, 13.0 ] [0.03, 0.11, 0.22, 0.07] 48.9
ABBAA [42.51, 13.52, 14.52, 11.17] [ 1.4, 31.6, 97.7, 10.7 ] [0.02, 0.11, 0.19, 0.06] 28.7
ABBAB [42.51, 13.71, 14.20, 11.21] [ 1.6, 158.1, 442.7, 7.1 ] [0.03, 0.25, 0.40, 0.05] 87.7
ABBBA [42.54, 13.71, 14.12, 11.20] [ 1.8, 121.1, 376.8, 7.5 ] [0.03, 0.22, 0.38, 0.05] 29.5
ABBBB [42.50, 13.62, 14.36, 11.20] [ 1.3, 8.0, 47.5, 10.0 ] [0.02, 0.06, 0.14, 0.06] 2.50
$\boldsymbol{\theta}^{true}=[\sigma_{y},A,n,a]^{true}=[42.51,13.63,14.35,11.19]$
@@
Table 4: Posterior summaries for 16 pre-determined load paths, each starting within an initial A load step. The parameter expected values, $\mathbb{E}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}$, marginal variances, $\mathbb{V}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}$ and 95% CI are reported for parameter vector $\boldsymbol{\theta}=[\sigma_{y},A,n,a]$ ($\sigma_{y}$ and $A$ reported in units of ksi) along with the generalized variance, which is the determinant of the posterior covariance. The 95% CI is represented as $\mathbb{E}_{\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi}}\,\pm\zeta$, where $\zeta$ is the reported value. For visual simplicity, horizontal load steps are indicated with an A and vertical load steps with a B. The best (AAAAA) and worst (ABABA) case scenarios as determined by parameter uncertainty are highlighted green and red, respectively. The material model parameters used for data generation were $\boldsymbol{\theta}^{true}=[42.51,13.63,14.35,11.19]$.
@@
In practice, not all 16 possible load paths would be evaluated experimentally, as the goal of this work is to identify the one preferred path based on its expected information content. Therefore, the load path was then chosen adaptively, as would be done in reality with the ICC framework, by calculating the EIG of the candidate load steps in $\Xi$ at each decision point as the cruciform was guided through the load path tree. Figure 9 illustrates the EIG for each candidate load step at each decision point. As the load steps progress, the EIG values steadily decrease, indicating that additional data from either load step has a waning influence on the calibration. In this study, an A load step was selected each time, resulting in a total load path of AAAAA. This choice was confirmed by the pre-determined load paths in Table 4, showing that AAAAA had the least collective parameter uncertainty given the data. While it was not guaranteed, this myopic decision-making led to the globally preferred full load path within the current tree structure. The selection of an AAAAA load path over alternatives, like ABABA, is not necessarily intuitive a priori. This outcome demonstrates that the BOED algorithm effectively identified the load path that yielded more informative data compared to other paths which might have been chosen based on subject-matter expertise.
Figure 9: The EIG was calculated for each candidate load step (A or B) at each decision point. Since the first load step was pre-determined to be A (pulling on the horizontal axis of the cruciform), the EIG was only calculated for steps 2-5. Load step A had the greater EIG at each load step, thus resulting in a total load path of AAAAA.
Fig. 10 shows a side-by-side comparison of the posterior contours from the AAAAA and ABABA calibrations so that a visual comparison can be made between the cases with least and greatest parameter uncertainty. Contours from each calibration are shown separately in Fig. 1 in B. The joint marginal posterior contours shown on the off-diagonal plots show the posterior distribution of two parameters simultaneously, which provide insights into their likely values based on the observed data and prior information. Each contour line represents a level of constant posterior density, with innermost contours indicating regions of higher probability and outer contours indicating lower probability for those parameter combinations. The contour shapes provide insight into the relationship between the parameters. Elliptical or elongated contours suggest a correlation between the two parameters, which is indicative of a trade-off relationship between the given parameters such that changing the parameters along the direction of correlation results a model output that is relatively stable (for a diffuse prior); circular contours imply independence. The spread of these distributions reflects uncertainty; wider distributions indicate greater variability and less confidence in the parameter estimates, whereas narrower distributions, with concentration around specific values, imply more precise estimates. The marginal posterior distribution of each parameter can be inferred by projecting the contours of the joint distributions onto the axes, providing a clearer understanding of individual parameter uncertainty. The marginal distributions are shown on the diagonals plots along with $\boldsymbol{\theta}^{true}$ (pink dashed line). $\boldsymbol{\theta}^{true}$ is also included with the joint marginal contours, indicated by a pink star.
The two load paths resulted in posterior distributions that exhibited different levels of joint and marginal parameter uncertainty as seen by the shape and spread of the contours, as well as different levels of parameter correlation as seen by the distinct amount of contour elongation (specific values of parameter correlation are reported in B). While the expected values align closely (Table 4), the marginal and joint posterior plots reveal subtle differences in the means when considering the small uncertainties in the parameters. This visual comparison elucidates the distinct outcomes of both calibrations; importantly, load path AAAAA resulted in lower parameter uncertainty compared to load path ABABA.
Figure 10: The posterior contours for load paths AAAAA and ABABA. Marginal posterior densities are plotted on the diagonals and joint-marginal densities on the off-diagonals. Marginal densities for the AAAAA calibration are plotted with a green line and for the ABABA calibration with a purple line. $\boldsymbol{\theta}^{true}$ is indicated with a dashed pink line and star.
The posterior probability is visualised over the surrogate model output by drawing 200 parameter samples from the posterior distribution $\boldsymbol{\theta}^{draw}\sim\pi(\boldsymbol{\theta}\>|\>\boldsymbol{y},\boldsymbol{\xi})$ and evaluating the surrogate model at each sample $\tilde{g}(\boldsymbol{\theta}^{draw})$. The corresponding posterior probability of the load values at load step 5 is shown in Fig. 11 for the AAAAA calibration. The model output at the MAP parameter estimate is shown with a dotted black line, and the 95% CI of the distribution over the load is shown with dashed black lines, indicating the interval in which the load values have a 95% probability of residing within.
The singular values obtained from $\tilde{g}(\boldsymbol{\theta}^{draw})$ were then used along with the 5 retained principal components to reconstruct the displacement field, $\text{disp}_{i}^{\text{recon}}=\text{disp}^{\text{PCA}}_{i}V^{*T}_{i}+\text{disp}^{\text{mean}}_{i},\,i\in[X,Y]$, so that there were 200 reconstructed instances of the displacement field. Fig. 12 shows the observed displacement field generated with $\boldsymbol{\theta}^{true}$ at the final load step of the AAAAA load path along with two line scans, across which the 95% CI was computed from the 200 reconstructed fields and plotted in Fig. 13. Given the relative scale of the displacements and the low uncertainty, the 95% CI interval appears as a single line. Overall, the high confidence in the parameter values translates to a high confidence in the resulting model output. The reader may refer to the Supplementary Material to view the posterior density over all QoIs and directional components for both the AAAAA and ABABA calibrations—global load values are shown in Fig. S8 and the displacements in Figs. S9-S12. Similar results were observed for the ABABA calibration: The posterior probability over the parameters translated to high confidence in the model output and recovered displacement field.
Figure 11: Posterior probability over the load for the AAAAA calibration for both the $X$ (left column) and $Y$ (right column) directional components. The load at the maximum a posteriori (MAP) parameter estimate is shown with a dotted black line, and the 95% credible interval (CI) of the distribution over the load is shown with dashed black lines, indicating the interval in which the load values have a 95% probability of residing within.
@@
Figure 12: The observed displacement field generated with $\boldsymbol{\theta}^{true}$ and added noise for the (a) X directional component and (b) Y directional component of the AAAAA load path at the final load step reported in microns. Lines (1) and (2) indicate the space over which the posterior probability of the displacement field is plotted in Fig. 13.
@@
Figure 13: Posterior probability over the nodal displacements for the AAAAA calibration across the two lines indicated in Fig. 12. The $X$ (left column) and $Y$ (right column) directional components are shown along Line 1 (top row) and Line 2 (bottom row). The 95% CI of the distribution over the displacement is shown with dashed black lines. Given the relative scale of the displacements and the low uncertainty, the 95% CI interval appears as a single line. The x-axis is the normalized location across each line scan.
@@ 4.3 Model Validation
Figure 14 presents the posterior predictive distribution over the load for the new ABBBA load path, derived from the AAAAA-calibrated model. The distribution, calculated from 200 samples, is organized by load step (rows) and directional component (columns). The figure also includes the observed data (solid red line), which was generated by running the model with the true parameters $\boldsymbol{\theta}^{true}$ for the ABBBA load path, incorporating noise to simulate real-world conditions. Additionally, the predicted load at the MAP parameter estimate is represented by the dotted black line, while the point-wise 95% prediction intervals (PI) are indicated by the dashed black lines. These intervals illustrate the uncertainty surrounding future observations based on the model’s predictions. Notably, for each load step and for both directional components, the observed data falls comfortably within the 95% PI, indicating a high probability under the assumed model. This alignment between the observed data and the predictive intervals provides compelling evidence that the model is well-calibrated, reinforcing its reliability for future predictions. Furthermore, the ability of the model to capture the underlying distribution of the load suggests that it effectively accounts for the inherent variability in the data.
@@
Figure 14: The AAAAA posterior predictive distribution for the new ABBBA load path is shown over the load for the $X$ (left column) and $Y$ (right column) directional components and at each load step. The observed data is shown with a solid red line, the load at the maximum a posteriori (MAP) parameter estimate is shown with a dotted black line, and the 95% prediction interval (PI) is shown with dashed black lines.
@@
Although it is challenging to visualize the posterior predictive distribution over the entire displacement field, two complementary visualizations are utilized here to gain insight on the predicted uncertainty. Of the 200 samples from the posterior predictive distribution, 3 were chosen to reconstruct corresponding displacement fields and are visualized in Fig. 15 for the $X$ directional component. The data collected at each load step in the ABBBA load path is shown along with 3 samples from the predictive distribution. This allows for a comparison to be made between the displacement field that is observed and that which is expected to be observed under the posterior distribution that was obtained. The multiple predictive samples serve to establish repeatability, or that the patterns seen in the displacement field did not merely happen by chance. There is no visible systematic discrepancy between the observed and predicted data, and it is reasonable to suggest that the observed data came from the same distribution as the predictions. The error between the validation data and the predictive fields is shown in Fig. 2 in B.1. This figure provides additional evidence of the repeatability of the results and confirms the absence of systematic discrepancies. Similar results were seen for the Y directional component in Figs. S15 and S17 in the Supplementary Material.
@@
Figure 15: Samples from the AAAAA posterior predictive distribution are shown for the $X$ directional component of the displacement field for the new ABBBA load path. The left column shows the observed data for each of the 5 load steps (LS) and the three right columns show three different predictive samples.
@@
The 95% PI of the distribution was calculated over the field point-wise from the 200 samples. Fig. 16 shows a full-field contour plot from which the 95% PI over Line 1 was extracted and plotted in Fig. 17 for all 5 load steps to show the predictive uncertainty. Plots in the left column show the 95% PI (dashed lines) along with the observed data (red line). Since distinguishing between the lines is difficult in some plots, the plots in the right column show the difference between the 95% PI and the observed data. The solid line represents the observations (centered at 0 difference), and the difference between the observations and the PIs are shown with dashed lines. Everywhere the observations fall outside the 95% PI coincides with locations where the upper bound difference is negative or the lower bound difference is positive; these regions are colored orange. Between 0% and 2.3% of the points along Line 1 have observations that fall outside the 95% PI for all load steps and both directional components, less than the expected amount of 5%. Similar observations were made for the predictive distribution over Line 2 in Fig. S20 in the Supplementary Material.
@@
Figure 16: The observed displacement field for the (a) X and (b) Y directional components of the ABBBA load path at the final load step reported in microns. Lines 1 marks the space over which the posterior predictive distribution is plotted in Fig. 17—the posterior predictive is shown over Line 2 in Fig. S20 in the Supplementary Material.
Figure 17: The AAAAA posterior predictive distribution is plotted over the Line 1 displacements (Fig. 16) for load path ABBBA. The two left columns show the 95% prediction interval (PI) of the distribution as well as the observations for the $X$ and $Y$ directional components for all 5 load steps (LS). The two right columns show the difference between the 95% PI and the observations. Regions where the observations fall outside the 95% PI are colored orange, and 0 is marked with a red line. The x-axis is the normalized location along Line 1.
Similar plots were made for the posterior predictive distribution from the ABABA calibration for the new ABBBA load path to serve as a visual comparison of the two cases and may be found in the Supplementary Material—Fig. S13 shows the posterior predictive density over the global load values and Figs. S14-S26 show the same over the displacement field for both the AAAAA— and ABABA—calibrated models. Additionally, a quantitative comparison is made between the predictive accuracy of the AAAAA vs. ABABA calibrations in C through the log predictive density (or log likelihood). The results show that the AAAAA load path generally yielded greater predictive probability than the ABABA load path.
In summary, the model validation showed that the model obtained through the calibration fit the data well. The posterior predictive distribution was visualized over the load at each load step as well as over the displacement field. There were no systematic discrepancies seen between the observed data and the predicted data. Additionally, the observed data looked plausible under the posterior predictive distribution; generally speaking, the observed data had high probability under the assumed model and fell within the 95% PI. Thus, the well-calibrated and validated model with uncertainty metrics can be used for decision-making with a quantified level of confidence. Importantly, the ICC framework achieved this efficiently and with greater credibility than would have been attained through alternative load paths.